OpenStudy (anonymous):

integral of x^2 * sqrt (2x +1)

OpenStudy (anonymous):

integration by parts with u = x^2 and dv = sqrt(2x + 1) are you familiar with it?

OpenStudy (wolfe8):

This is similar to the one I answered earlier. Do the same thing.

OpenStudy (anonymous):

btw you'll have to do it twice, the second time with u = x